# Unlocking the Hidden Potential of CLIP in Generalizable Deepfake Detection

### v1

@@ Abstract
This paper tackles the challenge of detecting partially manipulated facial deepfakes, which involve subtle alterations to specific facial features while retaining the overall context, posing a greater detection difficulty than fully synthetic faces. We leverage the Contrastive Language-Image Pre-training (CLIP) model, specifically its ViT-L/14 visual encoder, to develop a generalizable detection method that performs robustly across diverse datasets and unknown forgery techniques with minimal modifications to the original model. The proposed approach utilizes parameter-efficient fine-tuning (PEFT) techniques, such as LN-tuning, to adjust a small subset of the model’s parameters, preserving CLIP’s pre-trained knowledge and reducing overfitting. A tailored preprocessing pipeline optimizes the method for facial images, while regularization strategies, including L2 normalization and metric learning on a hyperspherical manifold, enhance generalization. Trained on the FaceForensics++ dataset and evaluated in a cross-dataset fashion on Celeb-DF-v2, DFDC, FFIW, and others, the proposed method achieves competitive detection accuracy comparable or outperforming much more complex state-of-the-art techniques. This work highlights the efficacy of CLIP’s visual encoder in facial deepfake detection and establishes a simple, powerful baseline for future research, advancing the field of generalizable deepfake detection. The code is available at: https://github.com/yermandy/deepfake-detection
@@ 3.1 Baseline model
As a baseline model, we use the CLIP-ViT-L/14 image encoder trained by OpenAI [1]. This model has $16\times 16+1$ tokens of dimension 1024. $16\times 16$ represents the number of visual patches and 1 refers to the classification token. We use only the classification token and discard all patch tokens. We freeze all weights to preserve the learned representations and train a single linear binary classifier with the cross-entropy loss function using the classification token as the input. In total, we train only weights $W\in\mathbb{R}^{1024\times 2}$ and biases $b\in\mathbb{R}^{2}$. To our knowledge, this baseline in the context of detection of AI-generated content was first introduced in [16] paper.
@@ 3.2 Dataset preprocessing
We found that while working with facial images, it is crucial to perform standardized facial processing for training and validation splits. We assume that when face swapping or face reenactment is applied to a facial image, the detector expects to see forgery cues in certain regions, which improves performance and makes the task easier to solve. Also, for face reenactment, where only a small area is changed (e.g. lip movements), the resolution plays an important role. For the standardized processing, we employ the pipeline proposed in DeepfakeBench [4], see further details in Section 4.2.
@@ 3.4 Regularization of the feature manifold to a hypersphere
We observed that restricting the features (outputs of the penultimate, before classification layer) to unit length by L2 normalization can improve the generalization to unseen datasets and forgery methods. This normalization can be viewed as a projection from Euclidean space to the hypersphere manifold. It was empirically observed that restricting features to live in a hypersphere helps to achieve better results in face-related tasks such as face recognition [26, 27]. From the perspective of a linear classifier, classification is akin to separating a spherical cap from the rest of a hypersphere. This encourages the model to create clusters on a hyperspherical cap and serves as a type of regularization.
@@ 3.6 Latent augmentations
Similarly to [15], we experiment with adding augmentations to extracted features. As we constrain features to live in a hyperspherical manifold, we take another approach to augmenting the latent space. We use spherical linear interpolation (slerp) between features from the same class in the same batch. For each feature $x_{i}$ in a batch, we randomly sample a paired feature $x_{j}$ having the same class $y_{i}=y_{j}$ and sample interpolation parameter $t\sim\mathcal{U}(0,1)$ from the uniform distribution. The slerp can then be defined as:
$\text{slerp}(\mathbf{x}_{i},\mathbf{x}_{j};t)=\sin((1-t)\theta)\,\mathbf{x}_{i}+\sin(t\theta)\,\mathbf{x}_{j}$ (1)
@@ 4.1 Datasets
For training and comparison purposes, we are using the FaceForensics++ (FF++) dataset [6] which consists of 3600 videos, from which 720 are real videos and the rest $4\times 720$ come from four different deepfake forgery methods, see Table 4.1. The dataset was released in three different compression levels. We employ the c23 version to facilitate a fair comparison with alternative methods.
We noticed that the usage of the FF++ validation set does not resemble the distribution of the test set and is very similar to the training set. The comparison of methods used in the ablation studies is not possible in these settings as each of them achieves the best scores quickly and does not show the overfitting effects. The FF++ validation set does not provide crucial information on how methods perform during longer training. Because of this, we create our own validation set consisting of a subset (15%) of the test set as well as samples from other deepfake datasets and with deepfakes created by us using FaceFusion software  [^2] https://github.com/facefusion/facefusion. In our experiments, we will show that using this validation set, we can find the set of regularization techniques that allows us to select most checkpoints without compromising on performance on the validation set.
@@
Table 1: Train dataset statistics. All models were trained on FF++ [6] dataset. The training split consists of 3600 videos, from which we sampled 22352 frames. There are 720 videos of the real class and $4\times 720$ videos belonging to four deepfake methods.
@@
Table 2: Test dataset statistics. Values represent the number of videos in real and fake classes used in our experiments. Values in parenthesis represent the number of videos missing from our test set compared to the original datasets due to the failure of the DLIB [29] object detector.
@@ 5 Results
Table 3: Generalization to unseen datasets and forgery methods trained on FF++ dataset. Reported values are video-level AUROC as described in Section 4.3. Results of other methods are taken from their original papers. Values with superscript citations are taken from the papers referenced in superscripts.

### v2

@@ Abstract
This paper tackles the challenge of detecting partially manipulated facial deepfakes, which involve subtle alterations to specific facial features while retaining the overall context, posing a greater detection difficulty than fully synthetic faces. We leverage the Contrastive Language-Image Pre-training (CLIP) model, specifically its ViT-L/14 visual encoder, to develop a generalizable detection method that performs robustly across diverse datasets and unknown forgery techniques with minimal modifications to the original model. The proposed approach utilizes parameter-efficient fine-tuning (PEFT) techniques, such as LN-tuning, to adjust a small subset of the model’s parameters, preserving CLIP’s pre-trained knowledge and reducing overfitting. A tailored preprocessing pipeline optimizes the method for facial images, while regularization strategies, including L2 normalization and metric learning on a hyperspherical manifold, enhance generalization. Trained on the FaceForensics++ dataset and evaluated in a cross-dataset fashion on Celeb-DF-v2, DFDC, FFIW, and others, the proposed method achieves competitive detection accuracy comparable to or outperforming much more complex state-of-the-art techniques. This work highlights the efficacy of CLIP’s visual encoder in facial deepfake detection and establishes a simple, powerful baseline for future research, advancing the field of generalizable deepfake detection. The code is available at: https://github.com/yermandy/deepfake-detection
@@ 3.1 Baseline model
As a baseline model, we use the CLIP-ViT-L/14 image encoder trained by OpenAI [1]. This model has $16\times 16+1$ tokens of dimension 1024. $16\times 16$ represents the number of visual patches, and 1 refers to the classification token. We use only the classification token and discard all patch tokens. We freeze all weights to preserve the learned representations and train a single linear binary classifier with the cross-entropy loss function using the classification token as the input. In total, we train only weights $W\in\mathbb{R}^{1024\times 2}$ and biases $b\in\mathbb{R}^{2}$. To our knowledge, this baseline in the context of detection of AI-generated content was first introduced in [16] paper.
@@ 3.2 Dataset preprocessing
We found that while working with facial images, it is crucial to perform standardized facial processing for training and validation splits. We assume that when face swapping or face reenactment is applied to a facial image, the detector expects to see forgery cues in certain regions, which improves performance and makes the task easier to solve. Also, for face reenactment, where only a small area is changed (e.g., lip movements), the resolution plays an important role. For the standardized processing, we employ the pipeline proposed in DeepfakeBench [4], see further details in Section 4.2.
@@ 3.4 Regularization of the feature manifold to a hypersphere
We observed that restricting the features (outputs of the penultimate layer, before the classification layer) to unit length by L2 normalization can improve the generalization to unseen datasets and forgery methods. This normalization can be viewed as a projection from Euclidean space to the hypersphere manifold. It was empirically observed that restricting features to live in a hypersphere helps to achieve better results in face-related tasks such as face recognition [26, 27]. From the perspective of a linear classifier, classification is akin to separating a spherical cap from the rest of a hypersphere. This encourages the model to create clusters on a hyperspherical cap and serves as a type of regularization.
@@ 3.6 Latent augmentations
Similarly to [15], we experiment with adding augmentations to extracted features. As we constrain features to live in a hyperspherical manifold, we take another approach to augmenting the latent space. We use spherical linear interpolation (slerp) between features from the same class in the same batch. For each feature $x_{i}$ in a batch, we randomly sample a paired feature $x_{j}$ having the same class $y_{i}=y_{j}$ and sample an interpolation parameter $t\sim\mathcal{U}(0,1)$ from the uniform distribution. The slerp can then be defined as:
$\text{slerp}(\mathbf{x}_{i},\mathbf{x}_{j};t)=\frac{\sin((1-t)\theta)}{\sin\theta}\mathbf{x}_{i}+\frac{\sin(t\theta)}{\sin\theta}\mathbf{x}_{j}$ (1)
@@ 4.1 Datasets
For training and comparison purposes, we are using the FaceForensics++ (FF++) dataset [6], which consists of 3600 videos, from which 720 are real videos and the rest $4\times 720$ come from four different deepfake forgery methods, see Table 4.1. The dataset was released in three different compression levels. We employ the c23 version to facilitate a fair comparison with alternative methods.
We noticed that the usage of the FF++ validation set does not resemble the distribution of the test set and is very similar to the training set. The comparison of methods used in the ablation studies is not possible in these settings, as each of them achieves the best scores quickly and does not show the overfitting effects. The FF++ validation set does not provide crucial information on how methods perform during longer training. Because of this, we create our own validation set consisting of a subset (15%) of the test set as well as samples from other deepfake datasets and with deepfakes created by us using FaceFusion software  [^2] https://github.com/facefusion/facefusion. In our experiments, we will show that using this validation set, we can find the set of regularization techniques that allows us to select most checkpoints without compromising on performance on the validation set.
@@
Table 1: Train dataset statistics. All models were trained on the FF++ [6] dataset. The training split consists of 3600 videos, from which we sampled 22352 frames. There are 720 videos of the real class and $4\times 720$ videos belonging to four deepfake methods.
@@
Table 2: Test dataset statistics. Values represent the number of videos in real and fake classes used in our experiments. Values in parentheses represent the number of videos missing from our test set compared to the original datasets due to the failure of the DLIB [29] object detector.
@@ 5 Results
Table 3: Generalization of models trained on the FF++ dataset to unseen datasets and forgery methods. Reported values are video-level AUROC as described in Section 4.3. Results of other methods are taken from their original papers. Values with superscript citations are taken from the papers referenced in superscripts.